# SynthSoM: A synthetic intelligent multi-modal sensing-communication dataset for Synesthesia of Machines (SoM)

### v1

@@ Methods
Figure 6 depicts the visualization and statistical properties of RF channels under different conditions. Figures 6(a) and (b) give path loss heatmaps in the urban crossroad scenario with sub-6 GHz and mmWave frequency bands, respectively. Since path loss is proportional to frequency, the path loss under mmWave frequency band, ranging from $82.1$ dBm to $157.7$ dBm, is higher than that with sub-6 GHz frequency band, ranging from $67.0$ dBm to $138.8$ dBm. This phenomenon is consistent with the measurement result [35]. Figures 6(c)–(e) show propagation paths in the mountain road scenario under SISO, MIMO, and massive MIMO, respectively. With an increase in the number of antennas, the number of propagation paths increases, leading to a larger spatial diversity gain. This phenomenon is in agreement with the measurement result [36]. Figure 6(f) compares space cross-correlation functions (SCCFs) in the mountain road scenario with sub-6 GHz and mmWave frequency bands under MIMO and massive MIMO conditions. The SCCF with sub-6 GHz frequency band is lower than that with mmWave frequency band, which is the consistent with the measurement result [37]. This is because that, compared to mmWave frequency band, there are more propagation paths in sub-6 GHz frequency band, resulting in larger channel spatial diversity and lower spatial correlation. Similarly, the SCCF under massive MIMO condition with more propagation paths is lower than that under MIMO condition.
@@
Figure 7: Visualization of RF sensory data under different intelligent agent densities. (a) mmWave radar point cloud under high intelligent agent density. (b) mmWave radar point cloud under medium intelligent agent density. (c) mmWave radar point cloud under low intelligent agent density. (d) Range-power map under high intelligent agent density. (e) Range-power map under medium intelligent agent density. (f) Range-power map under low intelligent agent density.
@@
Figure 7 gives the visualization of RF sensory data under different intelligent agent densities. For clarity, the suburban fork scenario is taken as an example for analysis. Figures 7(a)–(c) show mmWave radar point clouds under high, medium, and low intelligent agent densities at an intersection, respectively. Compared to the low and medium intelligent agent densities, the intelligent agent number within the FoV of mmWave radar is greater in the high intelligent agent density, leading to a higher number of detected mmWave radar points. Figures 7(d)–(f) correspond to the range-power maps of Figs. 7(a)–(c), respectively. As the intelligent agent number detected by the mmWave radar increases, the number of power spectrum density peaks increases, which is consistent with the measurement result [38]. Furthermore, the power spectrum density peak corresponding to the low-rise building approximately $32$ m away on the opposite side of the road decreases with the increase in the intelligent agent density. This is due to the increased blockage caused by the intelligent agent in the middle of the road.
@@
Figure 8: Visualization of non-RF sensory data under different conditions. (a) RGB image in urban overpass scenario on a sunny morning. (b) RGB image in urban overpass scenario on a sunny night. (c) RGB image in urban wide lane scenario on a rainy morning. (d) Depth map in urban wide lane scenario on a rainy morning. (e) LiDAR point cloud in urban wide lane scenario on a rainy morning. (f) RGB image in urban wide lane scenario on a snowy morning. (g) Depth map in urban wide lane scenario on a snowy morning. (h) LiDAR point cloud in urban wide lane scenario on a snowy morning.
@@
Figure 8 illustrates the visualization of non-RF sensory data under different conditions. Figures 8(a) and (b) show the RGB images in the urban overpass scenario on sunny morning and night, respectively. Certainly, the visibility of RGB image in the morning is higher than that in the night. Figures 8(c)–(h) present RGB images, depth maps, and LiDAR point clouds in the urban wide lane scenario on rainy and snowy morning, respectively. In rainy and snowy days, there are noise data in the RGB image and the depth map, which is consistent with the measurement result [39]. Furthermore, the rainy and snowy weather conditions also have a significant impact on LiDAR point clouds. Specifically, the laser attenuation resulted from random scattering medium leads to the increase in range uncertainty and loss of points far away from the LiDAR device. The huge backscattered power from random droplets further leads to the presence of random scattered points near the LiDAR device. The aforementioned phenomenon related to the LiDAR point cloud is also in agreement with the measurement result [30].
@@
Collecting communication and multi-modal sensory data in the real world requires various expensive measurement devices. These measurement devices further require spatial-temporal calibration under dynamic scenarios, which is of huge difficulty. As a result, it is expensive and difficult to collect communication and multi-modal sensory data under dynamic scenarios in the real world. To overcome this challenge, we conduct a measurement campaign at the campus of Peking University to collect communication and multi-modal sensory data under dynamic scenarios, as shown in Fig. 9(a). Although the measurement scenario is not specific to the vehicular scenario, the transceiver is mobile with dynamic pedestrians and static trees/buildings in the vicinity, as shown in Fig. 9(b). Therefore, we try our best to approximate the measurement scenario to a dynamic vehicular scenario [^1] Strictly speaking, the velocity of the transceiver in the vehicular scenario is higher than that in the conducted measurement campaign. Nonetheless, we are not principally interested in predicting the path loss outcome. Conversely, the ML experiment aims to demonstrate that the performance on the synthetic data can be transferred to the real world, which is similar to the performance transferability [41].. In the measurement campaign, we simultaneously utilize the sensor, communication equipment, and control device. The sensor includes an RGB-D camera, which collects RGB images with 1280$\times$720 resolution and depth maps with 848$\times$480 resolution. The communication equipment, which functions as the main RF transceiver, contains a pair of universal software radio peripheral (USRP) X410 devices and a pair of frequency converters connected to the USRP, a phased-array antenna connected to the frequency converter in transmitter (Tx) and a horn antenna connected to the frequency converters in receiver (Rx). As a result, the path loss between Tx and Rx can be collected, where the carrier frequency is $28$ GHz with $200$ MHz communication bandwidth. The control device consists of a pair of x86 industrial personal computers (IPCs) for controlling the transmission/reception of signals and storage of communication data, i.e., path loss, along with an ARM IPC for controlling the collection and storage of sensory data, i.e., RGB images and depth maps. By utilizing the sensor, communication equipment, and control device, we collect $882$ snapshots of aligned RGB images, depth maps, and path loss in the real world.
@@
Batch Size 8
Number of Hidden Layers 4
@@
Epoch 100
Learning Rate 0.001
@@
Based on the collected RGB image, depth map, and path loss in the real world and the developed simulation platform, we conduct the ML experiment for performance transferability testing, i.e., TSTR and TRTR, as shown in Fig. 10. For TSTR, based on an object detection network, i.e., You Only Look Once (YOLO)-v5, we utilize simulated RGB images and depth maps to extract the environmental feature and path loss, which are utilized for training and validation via multilayer perceptron (MLP). Furthermore, we evaluate the path loss prediction result of the MLP trained on the synthetic data using the testing set from the measurement data. For TRTR, we exploit the measurement data for training, validation, and testing to obtain the path loss prediction result via YOLO-v5 and MLP. For a fair comparison, the hyperparameters of MLPs for both TSTR and TRTR are consistent, as shown in Table 3. Meanwhile, the dataset is divided into the training set, validation set, and testing set in the proportion of $3:1:1$ for both TSTR and TRTR. The accuracy of path loss prediction can be computed by $A=\frac{\left|PL_{\mathrm{pr}}-PL_{\mathrm{gt}}\right|}{P_{\mathrm{gt}}}$, where $PL_{\mathrm{pr}}$ is the prediction result and $PL_{\mathrm{gt}}$ is the ground truth. The accuracies of path loss prediction via TSTR and TRTR are $87.35\%$ and $90.06\%$, respectively. Further comparing TSTR with TRTR in Fig. 11(a) and Fig. 11(b), path loss prediction accuracies are highly similar between synthetic data in the SynthSoM dataset and measurement data. Also, the TSTR performs worse than the TRTR, which is consistent with the result [41]. To be more specific, we observe a gap of less than 10% in accuracy between synthetic and measurement datasets, which is acceptable for a synthetic dataset, e.g., Table 6 [42] and Tables 4, 5 [43]. As a consequence, the comparison indicates that the performance on the synthetic data in the SynthSoM dataset can be transferred to the real world.
@@
Figure 11: Path loss prediction results of TSTR and TRTR. (a) Prediction accuracy at some specific snapshots. (b) Probability of prediction relative error.
@@
A git repository is publicly available on GitHub (https://github.com/ZiweiHuang96/SynthSoM/tree/main/code). In this repository, based on the SynthSoM dataset, several Python and MATLAB scripts for benchmarking, visualization, and data pre-processing are available.
@@ Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 62125101, Grant 62341101, and Grant 62371273; in part by the Taishan Scholars Program under Grant tsqn202312307; in part by the Young Elite Scientists Sponsorship Program by CAST under Grant YESS20230372; in part by the Shandong Natural Science Foundation under Grant ZR2023YQ058; in part by the New Cornerstone Science Foundation through the XPLORER PRIZE; in part by the Xiaomi Young Talents Program; in part by the open research fund of National Mobile Communications Research Laboratory, Southeast University (No. 2025D04); and in part by the China National Postdoctoral Program for Innovative Talents under Grant BX20240007. The authors would like to thank Mengyuan Lu, Xujing Jia, and Ye Chen for their help in the collection of communication data via Wireless InSite and Weibo Wen, Xuanyu Liu, Junliang Lu, and Beinan Su for their help in the collection of sensory data via AirSim and WaveFarer.

### v2

@@ Methods
Figure 6 depicts the visualization and statistical properties of RF channels under different conditions. Figures 6(a) and (b) give path loss heatmaps in the urban crossroad scenario with sub-6 GHz and mmWave frequency bands, respectively. Since path loss is proportional to frequency, the path loss under mmWave frequency band, ranging from $82.1$ dBm to $157.7$ dBm, is higher than that with sub-6 GHz frequency band, ranging from $67.0$ dBm to $138.8$ dBm. This phenomenon is consistent with the measurement result [36]. Figures 6(c)–(e) show propagation paths in the mountain road scenario under SISO, MIMO, and massive MIMO, respectively. With an increase in the number of antennas, the number of propagation paths increases, leading to a larger spatial diversity gain. This phenomenon is in agreement with the measurement result [37]. Figure 6(f) compares space cross-correlation functions (SCCFs) in the mountain road scenario with sub-6 GHz and mmWave frequency bands under MIMO and massive MIMO conditions. The SCCF with sub-6 GHz frequency band is lower than that with mmWave frequency band, which is the consistent with the measurement result [38]. This is because that, compared to mmWave frequency band, there are more propagation paths in sub-6 GHz frequency band, resulting in larger channel spatial diversity and lower spatial correlation. Similarly, the SCCF under massive MIMO condition with more propagation paths is lower than that under MIMO condition. Figures 6(g)–(i) demonstrate path loss heatmaps in the urban crossroad scenario on sunny, rainy, and sunny days, respectively. In comparison with sunny days, path loss is much smaller on rainy and snowy days attributed to the obvious rain and snow attenuation at the mmWave frequency band. This is consistent with the measurement result [39, 40].
@@
Figure 7: Visualization of RF sensory data under different intelligent agent densities. (a) mmWave radar point cloud under high intelligent agent density. (b) mmWave radar point cloud under medium intelligent agent density. (c) mmWave radar point cloud under low intelligent agent density. (d) Range-power map under high intelligent agent density. (e) Range-power map under medium intelligent agent density. (f) Range-power map under low intelligent agent density. (g) mmWave radar point cloud on sunny days. (h) mmWave radar point cloud on rainy days. (i) mmWave radar point cloud on snowy days.
@@
Figure 7 gives the visualization of RF sensory data under different intelligent agent densities. For clarity, the suburban fork scenario is taken as an example for analysis. Figures 7(a)–(c) show mmWave radar point clouds under high, medium, and low intelligent agent densities at an intersection, respectively. Compared to the low and medium intelligent agent densities, the intelligent agent number within the FoV of mmWave radar is greater in the high intelligent agent density, leading to a higher number of detected mmWave radar points. Figures 7(d)–(f) correspond to the range-power maps of Figs. 7(a)–(c), respectively. As the intelligent agent number detected by the mmWave radar increases, the number of power spectrum density peaks increases, which is consistent with the measurement result [41]. Furthermore, the power spectrum density peak corresponding to the low-rise building approximately $32$ m away on the opposite side of the road decreases with the increase in the intelligent agent density. This is due to the increased blockage caused by the intelligent agent in the middle of the road. Figures 7(g)–(i) depict mmWave radar point clouds on sunny, rainy, and snowy days, respectively. The scattering and obstruction of electromagnetic waves are minimal, leading to precise object detection and a longer detection range on sunny days. However, the scattering and absorption of mmWave radar signals by raindrops on rainy days reduce the maximum detection range of the mmWave radar [42]. Also, the scattering and absorption of mmWave radar signals by snowflakes are more pronounced on snowy days, thus further reducing the maximum detection range and the accuracy of object detection [43].
@@
Figure 8: Visualization of non-RF sensory data under different conditions. (a) RGB image on sunny days. (b) RGB image on rainy days. (c) RGB image on snowy days. (d) Depth map on sunny days. (e) Depth map on rainy days. (f) Depth map on snowy days. (g) LiDAR point cloud on sunny days. (h) LiDAR point cloud on rainy days. (i) LiDAR point cloud on snowy days. (j) RGB image at dawn. (k) RGB image in the morning. (l) RGB image at night.
@@
Figure 8 illustrates the visualization of non-RF sensory data in the mountain road scenario under different times of the day and weather conditions. Figures 8(a)–(c) show RGB images on sunny, rainy, and snowy days, respectively. Figures 8(d)–(f) show depth maps on sunny, rainy, and snowy days, respectively. On rainy and snowy days, there are noise data in the RGB image and the depth map, which is consistent with the measurement result [44]. Figures 8(g)–(i) show LiDAR point clouds on sunny, rainy, and snowy days, respectively. Also, the rainy and snowy weather conditions have a significant impact on LiDAR point clouds. Specifically, the laser attenuation resulted from random scattering medium leads to the increase in range uncertainty and loss of points far away from the LiDAR device. The huge backscattered power from random droplets further leads to the presence of random scattered points near the LiDAR device. The aforementioned phenomenon related to the LiDAR point cloud is also in agreement with the measurement result [30]. Finally, Figures 8(j)–(l) show RGB images at dawn, morning, and night, respectively. It is certain that the visibility of RGB image in the morning is higher than that of RGB images at dawn/night.
@@
Collecting communication and multi-modal sensory data in the real world requires various expensive measurement devices. These measurement devices further require spatial-temporal calibration under dynamic scenarios, which is of huge difficulty. As a result, it is expensive and difficult to collect communication and multi-modal sensory data under dynamic scenarios in the real world. To overcome this challenge, we conduct a measurement campaign at the campus of Peking University to collect communication and multi-modal sensory data under dynamic scenarios, as shown in Fig. 9(a). Although the measurement scenario is not specific to the vehicular scenario, the transceiver is mobile with dynamic pedestrians and static trees/buildings in the vicinity, as shown in Fig. 9(b). Therefore, we try our best to approximate the measurement scenario to a dynamic vehicular scenario. Strictly speaking, the velocity of the transceiver in the vehicular scenario is higher than that in the conducted measurement campaign. Nonetheless, we are not principally interested in predicting the path loss outcome. Conversely, the ML experiment aims to demonstrate that the performance on the synthetic data can be transferred to the real world, which is similar to the performance transferability [46]. In the measurement campaign, we simultaneously utilize the sensor, communication equipment, and control device. The sensor includes an RGB-D camera, which collects RGB images with 1280$\times$720 resolution and depth maps with 848$\times$480 resolution. The communication equipment, which functions as the main RF transceiver, contains a pair of universal software radio peripheral (USRP) X410 devices and a pair of frequency converters connected to the USRP, a phased-array antenna connected to the frequency converter in Tx and a horn antenna connected to the frequency converters in Rx. As a result, the path loss between Tx and Rx can be collected, where the carrier frequency is $28$ GHz with $200$ MHz communication bandwidth. The control device consists of a pair of x86 industrial personal computers (IPCs) for controlling the transmission/reception of signals and storage of communication data, i.e., path loss, along with an ARM IPC for controlling the collection and storage of sensory data, i.e., RGB images and depth maps. By utilizing the sensor, communication equipment, and control device, we collect $882$ snapshots of aligned RGB images, depth maps, and path loss in the real world.
@@
Batch Size 2
Number of Hidden Layers 3
@@
Epoch 40
Learning Rate 0.01
@@
Based on the collected RGB image, depth map, and path loss in the real world and the developed simulation platform, we conduct the ML experiment for performance transferability testing, i.e., TSTR and TRTR, as shown in Fig. 10. For TSTR, based on an object detection network, i.e., You Only Look Once (YOLO)-v5, we utilize simulated RGB images and depth maps to extract the environmental feature and path loss, which are utilized for training and validation via multilayer perceptron (MLP). Furthermore, we evaluate the path loss prediction result of the MLP trained on the synthetic data using the testing set from the measurement data. For TRTR, we exploit the measurement data for training, validation, and testing to obtain the path loss prediction result via YOLO-v5 and MLP. For a fair comparison, the hyperparameters of MLPs for both TSTR and TRTR are consistent, as shown in Table 3. Meanwhile, the dataset is divided into the training set, validation set, and testing set in the proportion of $3:1:1$ for both TSTR and TRTR. Therefore, the training set, validation set, and testing set contain 530 snapshots, 176 snapshots, and 176 snapshots of data, respectively.
The accuracy of path loss prediction can be computed by $A=\frac{\left|PL_{\mathrm{pr}}-PL_{\mathrm{gt}}\right|}{P_{\mathrm{gt}}}$, where $PL_{\mathrm{pr}}$ is the prediction result and $PL_{\mathrm{gt}}$ is the ground truth. Overall, the accuracies of path loss prediction via TSTR and TRTR are $89.28\%$ and $90.35\%$, respectively. Figure 11(a) depicts the prediction accuracies at all snapshots, i.e., 176 snapshots, of data in the testing set via TSTR and TRTR. Figure 11(b) further presents the probabilities of prediction relative error at all snapshots of data in the testing set via TSTR and TRTR. It can be seen from Figs. 11(a) and (b) that path loss prediction accuracies are highly similar between synthetic data in the SynthSoM dataset and measurement data. Also, the TSTR performs worse than the TRTR, which is consistent with the result [46]. To be more specific, we observe a gap of less than 10% in accuracy between synthetic and measurement datasets, which is acceptable for a synthetic dataset, e.g., Table 6 [47] and Tables 4, 5 [48]. As a consequence, the comparison indicates that the performance on the synthetic data in the SynthSoM dataset can be transferred to the real world. The corresponding code is publicly available on GitHub https://github.com/ZiweiHuang96/SynthSoM/tree/main/code.
@@
Figure 11: Path loss prediction results of TSTR and TRTR. (a) Prediction accuracy at all snapshots of data in the testing set. (b) Probability of prediction relative error at all snapshots of data in the testing set.
@@
A git repository is publicly available on GitHub https://github.com/ZiweiHuang96/SynthSoM/tree/main/code. In this repository, based on the SynthSoM dataset, several Python and MATLAB scripts for benchmarking, visualization, and data pre-processing are available.
@@ Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 62125101, Grant 62341101, and Grant 62371273; in part by the Taishan Scholars Program under Grant tsqn202312307; in part by the Young Elite Scientists Sponsorship Program by CAST under Grant YESS20230372; in part by the Shandong Natural Science Foundation under Grant ZR2023YQ058; in part by the New Cornerstone Science Foundation through the XPLORER PRIZE; in part by the Xiaomi Young Talents Program; in part by the open research fund of National Mobile Communications Research Laboratory, Southeast University (No. 2025D04); in part by the Beijing Natural Science Foundation under Grant 4254067; in part by the China National Postdoctoral Program for Innovative Talents under Grant BX20240007; and in part by the China Postdoctoral Science Foundation under Grant 2024M760111. The authors would like to thank Mengyuan Lu, Xujing Jia, and Ye Chen for their help in the collection of communication data via Wireless InSite and Weibo Wen, Xuanyu Liu, Junliang Lu, and Beinan Su for their help in the collection of sensory data via AirSim and WaveFarer.